ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

4. Algebraic Geometry

1991 MS Classification: Yu, Rupert Wei Tze, Université de Poitiers, France Lattice polytopes for certain Demazure modules of sl_{n+1} We present results on the existence of lattice convex polytopes associated to certain Demazure modules of sl_{n+1} in view of constructing degenerations of generalized Schubert varieties of SL(n + 1) into toric varieties. The class of Schubert varieties presented here is strictly bigger than the one obtained by N. Gonciulea and V. Lakshmibai, see Transformation Groups, volume 1, no. 3 (1996) 215-248. In particular, it includes all generalized Schubert varieties in sl_{3}. Applications to the theory of representation will be presented. These results are joint work with R. Dehy.